OpenStudy (anonymous):

Convert the polar coordinates to rectangular. (-3.5, 0.916)

OpenStudy (anonymous):

help?

zepdrix (zepdrix):

These coordinates are in polar currently?\[\Large\rm (r,~\theta)=(-3.5,~0.916)\]

OpenStudy (anonymous):

yes, and need converted to (x,y)

zepdrix (zepdrix):

Use these for converting from `Polar` to `Cartesian`, \[\Large\rm x=r \cos \theta, \qquad\qquad y=r \sin \theta\]Just use a calculator from there ^^

OpenStudy (anonymous):

thats what I did, but it's not accepting that answer

OpenStudy (anonymous):

wants exact values if possible...

OpenStudy (anonymous):

I get (-2.13, -2.78) when rounding to two decimal places

OpenStudy (anonymous):

which seems right, because it's in quadrant 3..